x = -3y - 1
x = -2y - 4

Answer by Cromlix(4381)

Hi there,
x = -3y - 1
x = -2y - 4
.....
x + 3y = -1
x + 2y = -4
Subtract bottom from top
y = 3
Substitute y = 3 into:
x = -3y - 1
x = -3(3) - 1
x = -9 - 1
x = -10
{-10,3}
Hope this helps:-)